Please help me with this problem. I really appreciate it.
Two angles are complementary. The sum of the measure of the first angle and one-fourth the second angle is 83.25
Find the measure of the smaller angle?
What is the measure of the other angle?

Answer by checkley77(12844)

x+y=90 or x=90-y
x+y/4=83.25
90-y+y/4=83.25
-y+y/4=83.25-90
-3y/4=-6.75
-3y=4*-6.75
y=-27/-3
y=9 is the smaller angle.
x+9=90
x=90-9
x=81 is the larger angle.
Proof:
81+9/4=83.25
81+2.25=83.25
83.25=83.25